OpenStudy (anonymous):

Solve for x. log6(2x + 3) = 3

OpenStudy (mertsj):

\[6^3=2x+3\]

OpenStudy (anonymous):

log6(2x + 3) = 3 6(2x + 3) = 10^3 2x+3 = 1000/6 x= 491/6

OpenStudy (anonymous):

\[\log_b(x)=y\iff b^y=x\] you need to know who to switch back and forth easily , or rather you need to be able to switch easily for your math class

Directrix (directrix):

6^3 = 2x + 3 by definition of logs 216 = 2x + 3 213 = 2x 106.5 = x

OpenStudy (anonymous):

you get to mertsj answer in one step by rewriting in exponential form

OpenStudy (mertsj):

Listen to Satellite. He is the wise one.

OpenStudy (anonymous):

I have just found that 6 is the base. Sorry for that!